x+y=30
3x+4y=100
Solve for x&y
There would be___ 3-point questions and ___ 4-point questions
I need lots of detail on how to do this, please.
:)

Answer by math_tutor2020(3820)


Solve the first equation for y
x+y = 30
y = 30-x

Then plug this into the second equation
3x+4y = 100
3x+4( y )= 100
3x+4( 30-x )= 100

At this point we solve for x
3x+4( 30-x )= 100
3x+120-4x= 100
-x+120= 100
-x= 100-120
-x = -20
x = 20

Use this to find y
y = 30-x
y = 30-20
y = 10

Summary:
x = 20 and y = 10
The two lines cross at (20,10).